# RICE: Breaking Through the Training Bottlenecks of Reinforcement Learning with Explanation

### v1

@@ 5 Discussion and Conclusion
We believe that RICE provides a versatile framework for refining a pre-trained DRL agent. Though we mainly investigate the case when the target agent is trained by the PPO algorithm, our framework can also apply to the case when the agent is trained by other algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018). In order to preserve the theoretical guarantee, we can use imitation learning techniques such as Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) to learn an approximated policy network and use RICE to further refine it. Figure 3 demonstrates the advantage of our refining method against other baselines. We provide additional detailed experiment results in Appendix E.
RICE is suitable for DRL applications that are trained within controllable environment (e.g., simulators), in order to generate explanations. In fact, most of today’s DRL applications rely on some form of simulator for their training. For example, for safety-critical applications such as autonomous driving, DRL agents are usually designed, trained, and tested in a simulated environment first before moving them to real-world testing. Simulation platforms broadly include Carla (Dosovitskiy et al., 2017) and MetaDrive (Li et al., 2022) which have been used to facilitate the training of DRL agents (Zhang et al., 2021; Wang et al., 2023; Peng et al., 2022). Therefore, RICE should be applicable to such DRL systems (especially during their development phase) for refining a pre-trained DRL agent.
@@ Acknowledgements
This project was supported in part by NSF Grants 2225234, 2225225, 2229876, 1955719, and 2055233.
@@
This paper presents work whose goal is to advance the field of reinforcement learning with explanation. There are many potential social impacts of our work. Our approach provides a feasible solution to break through the training bottlenecks of reinforcement learning with explanation, which is an automatic process and saves manual effort. However, it is also worth noting the potential negative societal impacts of our work, our approach to refine RL agents may lead to the magnification of adversarial or biased RL agents. Nevertheless, our explanation method also has the capability of understanding the decision-making process of adversarial or biased RL agents, which enables the mitigation of their harmful effects.
@@
$\eta(\bar{\pi})-\eta(\pi)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}A^{\pi}\left(s,\bar{\pi}(s)\right),$ (4)
@@
Note that the perturbed policy $\bar{\pi}$ is a mixture of the target agent’s policy $\pi$ and a random policy $\pi^{r}$ (i.e., $\bar{\pi}(\cdot|s)=\tilde{\pi}(a^{e}=0|s)\pi(\cdot|s)+\tilde{\pi}(a^{e}=1|s)\pi^{r}(\cdot|s)$). Denote the probability of the mask network outputting 0 at state $s$ as $\tilde{\pi}(a^{e}=0|s)=\xi(s)$ and the probability of the mask network outputting 1 at state $s$ as $\tilde{\pi}(a^{e}=1|s)=1-\xi(s)$ Given the fact that $A^{\pi}\left(s,\pi(s)\right)=\mathbb{E}_{a\sim\pi(s)}A^{\pi}(s,a)=0$, we have
$\begin{split}\eta(\bar{\pi})-\eta(\pi)&=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}A^{\pi}\left(s,\bar{\pi}(s)\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\mathbb{E}_{a\sim\bar{\pi}(s)}A^{\pi}\left(s,\bar{\pi}(s)\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\sum_{a}\bar{\pi}(s)A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\sum_{a}\xi(s)\pi(s)A^{\pi}\left(s,a\right)+\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\sum_{a}(1-\xi(s)\pi^{r}(s)A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\xi(s)\mathbb{E}_{a\sim\pi(s)}A^{\pi}\left(s,a\right)+\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}(1-\xi(s))\mathbb{E}_{a\sim\pi^{r}(s)}A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}(1-\xi(s))\mathbb{E}_{a\sim\pi^{r}(s)}A^{\pi}\left(s,a\right)\leq 0.\\ \end{split}$ (5)
@@ Appendix B Theoretical Guarantee
Since our explanation method provides the importance of each state, we could view the sampling based on the state’s importance as a reweighting of the state occupancy measure. Mathematically, it can be expressed as $d_{\rho}^{\hat{\pi}}(s)=d_{\rho}^{\pi}(s)w(s)$, where $\hat{\pi}$ is the equivalent policy of reweighting the original policy $\pi$ and $w(s)$ is the weight provided by the mask network. Although the mask network takes the current input state as input, the state transition is implicitly associated with the related action at these states. As such, we have $d_{\rho}^{\hat{\pi}}(s,a)=d_{\rho}^{\pi}(s,a)w(s,a)$.
@@
Suppose two policies $\hat{\pi}$ and $\pi$ satisfy $g\left(\hat{\pi}(a\mid s)\right)=g(\pi(a\mid s))+h\left(s,Q^{\pi}(s,a)\right)$ Where $g(.)$ is a monotonically increasing function, and $h(s,\cdot)$ is monotonically increasing for any fixed $s$ Then we have $V^{\hat{\pi}}(s)\geq V^{\pi}(s)$, $\forall s\in\mathcal{S}$.
@@
$\mathcal{A}_{1}\triangleq\left\{a\in\mathcal{A}\mid\hat{\pi}(a\mid s)\geqslant\pi(a\mid s)\right\}.$
$\mathcal{A}_{2}\triangleq\left\{a\in\mathcal{A}\mid\hat{\pi}(a\mid s)<\pi(a\mid s)\right\}.$
@@
$\begin{split}h\left(s,Q^{\pi}(s,a_{1})\right)&=g\left(\hat{\pi}(a_{1}\mid s)\right)-g(\pi(a_{1}\mid s))\\ &\geq 0\\ &\geq g\left(\hat{\pi}(a_{2}\mid s)\right)-g(\pi(a_{2}\mid s))\\ &=h\left(s,Q^{\pi}(s,a_{2})\right).\end{split}$ (7)
@@
$\begin{split}&\sum_{a\in\mathcal{A}}\hat{\pi}(a\mid s)Q^{\pi}(s,a)-\sum_{a\in A}\pi(a\mid s)Q^{\pi}(s,a)\\ =&\sum_{a_{1}\in\mathcal{A}_{1}}\left(\hat{\pi}\left(a_{1}\mid s\right)-\pi(a_{1}\mid s)\right)Q^{\pi}\left(s,a_{1}\right)\\ &+\sum_{a_{2}\in\mathcal{A}_{2}}\left(\hat{\pi}\left(a_{2}\mid s\right)-\pi(a_{2}\mid s)\right)Q^{\pi}(s,a_{2})\\ \geq&\sum_{a_{1}\in\mathcal{A}_{1}}\left(\hat{\pi}\left(a_{1}\mid s\right)-\pi\left(a_{1}\mid s\right)\right)q(s)\\ &+\sum_{a_{2}\in\mathcal{A}_{2}}\left(\hat{\pi}\left(a_{2}\mid s\right)-\pi\left(a_{2}\mid s\right)\right)q(s)\\ =&q(s)\sum_{a\in\mathcal{A}}\left(\pi^{\prime}(a\mid s)-\pi(a\mid s)\right)\\ =&0.\\ \end{split}$ (8)
Let $V_{0}(s)=V^{\pi}(s)$. And we denote the value function of following $\hat{\pi}$ for $l$ steps then following $\pi$ as $V_{l}(s)=E_{a\sim\hat{\pi}(.|s)}\left[E_{s^{\prime},r\mid s,a}\left(r+\gamma V_{l-1}\left(s^{\prime}\right)\right)\right]$ if $l\geq 1$.
@@
$V_{1}(s)$ $=E_{a\sim\hat{\pi}(.|s)}\left[E_{s^{\prime},r\mid s,a}\left(r+\gamma V^{\pi}(s^{\prime})\right)]\right.$ (9)
@@
$V_{l+1}(s)=E_{a\sim\hat{\pi}}\left[E_{s^{\prime},r\mid s,a}\left(r+V_{l}\left(s^{\prime}\right)\right)\right],$
$V_{l}(s)=E_{a\sim\hat{\pi}}\left[E_{s^{\prime},r\mid s,a}\left(r+V_{l-1}\left(s^{\prime}\right)\right)\right],$
@@
Based on the Proposition B.2, if we choose $g$ as a logarithmic function and $h=log(w(s,a))+log(d_{\rho}^{\pi}(s))-log(d_{\rho}^{\hat{\pi}}(s))$, we can easily verify that our MaskNet-based sampling approach is equivalent to sampling from a better policy $\hat{\pi}$.
@@ B.3 Proof of Theorem 3
$\begin{split}\epsilon&>\sum_{s\in\mathcal{S}}d_{\mu}^{\pi^{\prime}}(s)\left[\max_{a}A^{\pi^{\prime}}(s,a)\right]\\ &>\sum_{s\in\mathcal{S}}d_{\mu}^{\pi^{\prime}}(s)\left[\max_{a}A^{\pi^{\prime}}(s,a)\right]\\ &\geq\min_{s}\left(\frac{d_{\mu}^{\pi^{\prime}}(s)}{d_{\rho}^{\pi^{*}}(s)}\right)\sum_{s}d_{\rho}^{\pi^{*}}(s)\max_{a}A^{\pi^{\prime}}(s,a)\\ &\geq\left\|\frac{d^{\pi^{*}}_{\rho}}{d_{\mu}^{\pi^{\prime}}}\right\|_{\infty}^{-1}\sum_{s,a}d_{\rho}^{\pi^{*}}(s)\pi^{*}(a\mid s)A^{\pi^{\prime}}(s,a).\\ \end{split}$ (10)
@@
$V^{\pi^{*}}\left(\rho\right)-V^{\pi^{\prime}}\left(\rho\right)$ (11)
$=$ $\frac{1}{1-\gamma}E_{s\sim d_{\rho}^{\pi*}}E_{a\sim\pi^{*}(.\mid s)}\left[A^{\pi^{\prime}}(s,a)\right].$
@@ C.3 Additional Experiment Results
Hyper-parameter Choices in Experiment I-IV.Table 4 summarizes our hyper-parameter choices in Experiment I-IV. For all applications, we choose the coefficient of the intrinsic reward for training the mask network $\alpha$ as 0.01. The hyper-parameters $p$ and $\lambda$ for our refining method vary by application.
Table 4: Hyper-parameter choices in Experiment I-IV. “Selfish” represents Selfish Mining. “Cage” represents Cage Challenge 2. “Auto” represents Autonomous Driving. “Malware” represents Malware Mutation.
  $p$ 0.25 0.25 0.50 0.50 0.25 0.50 0.25 0.50
$\lambda$ 0.001 0.01 0.001 0.01 0.001 0.01 0.01 0.01
$\alpha$ 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
Sensitivity of Hyper-parameters $p$ and $\lambda$.   We provide the sensitivity results of $p$ in all applications in Figure 6. We observe that generally a mixture probability pf $p=0.25$ or $p=0.5$ is a good choice. Additionally, recall that we need to use the hyper-parameter $\lambda$ to balance the scale of the “true” environment reward and the exploration bonus. We test the sensitivity of $\lambda$ from the space {0.1, 0.01, 0.001}. Figure 7 reports the agent’s performance after refining under different settings of $\lambda$. We observe that our retaining method is insensitive to the choice of $\lambda$. The agent’s performance does not vary a lot with different settings of $\lambda$. But $\lambda=0.01$ gives the best performance in all applications except selfish mining.

### v3

@@ Acknowledgements
This project was supported in part by Northwestern University TGS Fellowship and NSF Grants 2225234, 2225225, 2229876, 1955719, and 2055233.
@@
This paper presents work whose goal is to advance the field of reinforcement learning with explanation. There are many potential social impacts of our work. Our approach provides a feasible solution to break through the training bottlenecks of reinforcement learning with explanation, which is an automatic process and saves manual effort.
However, it is also worth noting the potential negative societal impacts of our work. Some of the real-world applications we select such as malware mutation can create attack examples that may bring additional ethical concerns. In the realm of security research, the ultimate goal of these tasks is to generate stronger testing cases to enhance the defense, and it is standard practice. Take malware mutation as an example, the produced samples can be used to proactively improve the robustness and effectiveness of malware detection systems (e.g., through adversarial training), thereby benefiting cybersecurity defense (Yang et al., 2017).
@@
$\eta(\bar{\pi})-\eta(\pi)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\mathbb{E}_{a\sim\bar{\pi}(\cdot|s)}A^{\pi}\left(s,a\right),$ (4)
@@
Note that the perturbed policy $\bar{\pi}$ is a mixture of the target agent’s policy $\pi$ and a random policy $\pi^{r}$ (i.e., $\bar{\pi}(\cdot|s)=\tilde{\pi}(a^{e}=0|s)\pi(\cdot|s)+\tilde{\pi}(a^{e}=1|s)\pi^{r}(\cdot|s)$). Denote the probability of the mask network outputting 0 at state $s$ as $\tilde{\pi}(a^{e}=0|s)=\xi(s)$ and the probability of the mask network outputting 1 at state $s$ as $\tilde{\pi}(a^{e}=1|s)=1-\xi(s)$ Given the fact that $A^{\pi}\left(s,\pi(\cdot|s)\right)=\mathbb{E}_{a\sim\pi(s)}A^{\pi}(s,a)=0$, we have
$\begin{split}\eta(\bar{\pi})-\eta(\pi)&=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\mathbb{E}_{a\sim\bar{\pi}(\cdot|s)}A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\sum_{a}\bar{\pi}(a|s)A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\sum_{a}\xi(s)\pi(a|s)A^{\pi}\left(s,a\right)+\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\sum_{a}(1-\xi(s))\pi^{r}(a|s)A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}\xi(s)\mathbb{E}_{a\sim\pi(\cdot|s)}A^{\pi}\left(s,a\right)+\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}(1-\xi(s))\mathbb{E}_{a\sim\pi^{r}(\cdot|s)}A^{\pi}\left(s,a\right)\\ &=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\rho}^{\bar{\pi}}}(1-\xi(s))\mathbb{E}_{a\sim\pi^{r}(\cdot|s)}A^{\pi}\left(s,a\right)\leq 0.\\ \end{split}$ (5)
@@ Appendix B Theoretical Guarantee
Since our explanation method provides the importance of each state, we could view the sampling based on the state’s importance as a reweighting of the state occupancy measure. Mathematically, it can be expressed as $d_{\rho}^{\hat{\pi}}(s)=d_{\rho}^{\pi}(s)w(s)$, where $\hat{\pi}$ is the equivalent policy of reweighting the original policy $\pi$ and $w(s)$ is the weight provided by the mask network. Although the mask network takes the current input state as input, it implicitly considers the current action as well, as detailed by StateMask (Cheng et al., 2023). Consequently, a more accurate formulation is $d_{\rho}^{\hat{\pi}}(s,a)=d_{\rho}^{\pi}(s,a)w(s,a)$, where $w(s,a)$ represent the weight assigned by mask network.
@@
Suppose two policies $\hat{\pi}$ and $\pi$ satisfy $g\left(\hat{\pi}(a|s)\right)=g(\pi(a|s))+h\left(s,Q^{\pi}(s,a)\right)$ where $g(\cdot)$ is a monotonically increasing function, and $h(s,\cdot)$ is monotonically increasing for any fixed $s$ Then we have $V^{\hat{\pi}}(s)\geq V^{\pi}(s)$, $\forall s\in\mathcal{S}$.
@@
$\mathcal{A}_{1}\triangleq\left\{a\in\mathcal{A}|\hat{\pi}(a|s)\geqslant\pi(a|s)\right\}.$
$\mathcal{A}_{2}\triangleq\left\{a\in\mathcal{A}|\hat{\pi}(a|s)<\pi(a|s)\right\}.$
@@
$\begin{split}h\left(s,Q^{\pi}(s,a_{1})\right)&=g\left(\hat{\pi}(a_{1}|s)\right)-g(\pi(a_{1}|s))\\ &\geq 0\\ &\geq g\left(\hat{\pi}(a_{2}|s)\right)-g(\pi(a_{2}|s))\\ &=h\left(s,Q^{\pi}(s,a_{2})\right).\end{split}$ (7)
@@
$\begin{split}&\sum_{a\in\mathcal{A}}\hat{\pi}(a|s)Q^{\pi}(s,a)-\sum_{a\in A}\pi(a|s)Q^{\pi}(s,a)\\ =&\sum_{a_{1}\in\mathcal{A}_{1}}\left(\hat{\pi}\left(a_{1}|s\right)-\pi(a_{1}|s)\right)Q^{\pi}\left(s,a_{1}\right)+\sum_{a_{2}\in\mathcal{A}_{2}}\left(\hat{\pi}\left(a_{2}|s\right)-\pi(a_{2}|s)\right)Q^{\pi}(s,a_{2})\\ \geq&\sum_{a_{1}\in\mathcal{A}_{1}}\left(\hat{\pi}\left(a_{1}|s\right)-\pi\left(a_{1}|s\right)\right)q(s)+\sum_{a_{2}\in\mathcal{A}_{2}}\left(\hat{\pi}\left(a_{2}|s\right)-\pi\left(a_{2}|s\right)\right)q(s)\\ =&\hskip 5.0ptq(s)\sum_{a\in\mathcal{A}}\left(\pi^{\prime}(a|s)-\pi(a|s)\right)\\ =&\hskip 5.0pt0.\\ \end{split}$ (8)
Let $V_{0}(s)=V^{\pi}(s)$. And we denote the value function of following $\hat{\pi}$ for $l$ steps then following $\pi$ as $V_{l}(s)=E_{a\sim\hat{\pi}(.|s)}\left[E_{s^{\prime},r|s,a}\left(r+\gamma V_{l-1}\left(s^{\prime}\right)\right)\right]$ if $l\geq 1$.
@@
$V_{1}(s)$ $=E_{a\sim\hat{\pi}(.|s)}\left[E_{s^{\prime},r|s,a}\left(r+\gamma V^{\pi}(s^{\prime})\right)]\right.$ (9)
@@
$V_{l+1}(s)=E_{a\sim\hat{\pi}}\left[E_{s^{\prime},r|s,a}\left(r+V_{l}\left(s^{\prime}\right)\right)\right],$
$V_{l}(s)=E_{a\sim\hat{\pi}}\left[E_{s^{\prime},r|s,a}\left(r+V_{l-1}\left(s^{\prime}\right)\right)\right],$
@@
Based on the Proposition B.2, if we choose $g$ as a logarithmic function and $h=\log(w(s,a))+\log(d_{\rho}^{\pi}(s))-\log(d_{\rho}^{\hat{\pi}}(s))$, we can easily verify that our MaskNet-based sampling approach is equivalent to sampling from a better policy $\hat{\pi}$.
@@ B.3 Proof of Theorem 3
$\begin{split}\epsilon&>\sum_{s\in\mathcal{S}}d_{\mu}^{\pi^{\prime}}(s)\left[\max_{a}A^{\pi^{\prime}}(s,a)\right]\\ &\geq\min_{s}\left(\frac{d_{\mu}^{\pi^{\prime}}(s)}{d_{\rho}^{\pi^{*}}(s)}\right)\sum_{s}d_{\rho}^{\pi^{*}}(s)\max_{a}A^{\pi^{\prime}}(s,a)\\ &\geq\left\|\frac{d^{\pi^{*}}_{\rho}}{d_{\mu}^{\pi^{\prime}}}\right\|_{\infty}^{-1}\sum_{s,a}d_{\rho}^{\pi^{*}}(s)\pi^{*}(a|s)A^{\pi^{\prime}}(s,a).\\ \end{split}$ (10)
@@
$V^{\pi^{*}}\left(\rho\right)-V^{\pi^{\prime}}\left(\rho\right)=\frac{1}{1-\gamma}E_{s\sim d_{\rho}^{\pi*}}E_{a\sim\pi^{*}(.|s)}\left[A^{\pi^{\prime}}(s,a)\right].$ (11)
@@ C.1 Implementation Details
We implement the environment reset function similar to Ecoffet et al. (2019) to restore the environment to selected critical states. This method is feasible in our case, as we operate within simulator-based environments. However, in the real world, it may not be always possible to return to a certain state with the same sequences of actions due to the stochastic nature of state transition. It’s important to note that our framework is designed to be versatile and is indeed compatible with a goal/state-conditioned policy approach such as Ecoffet et al. (2021). Given a trajectory with an identified most important state, we can select the most important state as the final goal and select the en-route intermediate states as sub-goals. Then we can train an agent to reach the final goal by augmenting each state with the next goal and giving a goal-conditioned reward once the next goal is reached until all goals are achieved.
@@ C.3 Additional Experiment Results
Hyper-parameter Choices in Experiment I-V.Table 3 summarizes our hyper-parameter choices in Experiment I-V. For all applications, we choose the coefficient of the intrinsic reward for training the mask network $\alpha$ as 0.01. The hyper-parameters $p$ and $\lambda$ for our refining method vary by application.
Table 3: Hyper-parameter choices in Experiment I-V. “Selfish” represents Selfish Mining. “Cage” represents Cage Challenge 2. “Auto” represents Autonomous Driving. “Malware” represents Malware Mutation.
  $p$     0.25 0.25 0.50 0.50 0.25 0.50 0.25 0.50
$\lambda$     0.001 0.01 0.001 0.01 0.001 0.01 0.01 0.01
$\alpha$     0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
@@
Comparison with Self-Imitation Learning.   We compare RICE against the self-imitation learning (SIL) approach (Oh et al., 2018) across four MuJoCo games. We present the results presented in Table 5. These experiment results demonstrate that RICE consistently outperforms the self-imitation learning method. While self-imitation learning has the advantage of encouraging the agent to imitate past successful experiences by prioritizing them in the replay buffer, it cannot address scenarios where the agent (and its past experience) has errors or sub-optimal actions. In contrast, RICE constructs a mixed initial distribution based on the identified critical states (using explanation methods) and encourages the agent to explore the new initial states. This helps the agent escape from local minima and break through the training bottlenecks.
Table 5: Performance comparison between Self-Imitation Learning (SIL) and RICE on four MuJoCo tasks.
  Method     Hopper Walker2d Reacher HalfCheetah
  SIL     3646.46 (23.12) 3967.66 (1.53) -2.87 (0.09) 2069.80 (3.44)
Ours     3663.91 (20.98) 3982.79 (3.15) -2.66 (0.03) 2138.89 (3.22)
Impact of Other Explanation Methods.   We investigate the impact of other explanation methods (i.e., AIRS (Yu et al., 2023) and Integrated Gradients (Sundararajan et al., 2017)) on four Mujoco games. we fix the refining method and use different explanation methods to identify critical steps for refinement. The results are reported in Table 6. We observe that using the explanation generated by our mask network, the refining achieves the best outcome across all four applications. Using other explanation methods (Integrated Gradients and AIRS), our framework still achieves better results than the random baseline, suggesting that our framework can work with different explanation method choices.
Table 6: Performance comparison when using different explanation methods across four MuJoCo tasks.
  Task     Random Explanation Integrated Gradients AIRS Ours
  Hopper     3648.98 (39.06) 3653.24 (14.23) 3654.49 (8.12) 3663.91 (20.98)
Walker2d     3969.64 (6.38) 3972.15 (4.77) 3976.35 (2.40) 3982.79 (3.15)
Reacher     -3.11 (0.42) -2.99 (0.31) -2.89 (0.19) -2.66 (0.03)
HalfCheetah     2132.01 (0.76) 2132.81 (0.83) 2133.98 (2.52) 2138.89 (3.22)
Sensitivty of $p$ and $\lambda$ in Hopper game with an imitated PPO agent.   We report the sensitivity of hyper-parameters $p$ and $\lambda$ in Hopper game with an imitated PPO agent in Figure 6. We observe that in general, a mixture probability of $p=0.25$ or $p=0.5$ is a better choice. An RND bonus can facilitate the agent with faster refinement.
Sensitivity of Hyper-parameters $p$ and $\lambda$.   We provide the sensitivity results of $p$ in all applications in Figure 7. We observe that generally a mixture probability of $p=0.25$ or $p=0.5$ is a good choice. Additionally, recall that we need to use the hyper-parameter $\lambda$ to balance the scale of the “true” environment reward and the exploration bonus. We test the sensitivity of $\lambda$ from the space {0.1, 0.01, 0.001}. Figure 8 reports the agent’s performance after refining under different settings of $\lambda$. We observe that our retaining method is insensitive to the choice of $\lambda$. The agent’s performance does not vary a lot with different settings of $\lambda$. But $\lambda=0.01$ gives the best performance in all applications except selfish mining.